OpenStudy (anonymous):

People with type O-negative blood are universal donors. That is, any patient can receive a transfusion of O-negative blood. Only 8.5% of the American population have O-negative blood. If 12 people appear at random to give blood, what is the probability (±0.0001) that at least one of them is a universal donor?

OpenStudy (chosenmatt):

10 people not having O blood. C(10, 0)*(.072)^0*(1-.072)^(10-0) = 1*1*(.4736) = .4736. Almost done. Now 1 - .4736 = .526